OpenStudy (anonymous):

Need a little help--Instantaneous velocity: (derivatives) s=8t(squared)-10t+6; t=5. any help?,Thanks

OpenStudy (anonymous):

Just use the sum rule. The derivative of a sum is the sum of the derivatives of each term.

myininaya (myininaya):

s'=2*8t-10 s'(5)=16(5)-10

OpenStudy (anonymous):

so s=156?

OpenStudy (anonymous):

No, s'(5) = 156

myininaya (myininaya):

s'(5)=80-10=70 did i miss something?

OpenStudy (anonymous):

yeah, it's 70

OpenStudy (anonymous):

I didn't check his arithmetic. In anycase, this is the trouble with just solving people's problems myininaya. They don't understand what you did, so they don't understand the answer, or how to do it again next time.

OpenStudy (anonymous):

can you take it step by step for me ? now I am missing something , Thanks

myininaya (myininaya):

do you know how to find the derivative?

OpenStudy (anonymous):

yes

myininaya (myininaya):

what is s' = then?

OpenStudy (anonymous):

Prime

myininaya (myininaya):

what is s'=______________?

OpenStudy (anonymous):

oh - i got 156

myininaya (myininaya):

the derivative of 8t^2 is? the derivative of 10t is? the derivative of constant is?

myininaya (myininaya):

no clown i'm asking what is s' not what s' at some number is

OpenStudy (anonymous):

16t,10,0

myininaya (myininaya):

so s'=16t+10+0=16t+10

myininaya (myininaya):

oops s'=16t-10 i forgot the minus in the problem

myininaya (myininaya):

so what is s' at t=5

OpenStudy (anonymous):

Man your rough! but you make me think! ---- Thanks

myininaya (myininaya):

i'm rough lol

myininaya (myininaya):

do you get 70 now?

OpenStudy (anonymous):

Hold on Im old!

OpenStudy (anonymous):

yes, Thank you , Wow ! LOL

myininaya (myininaya):

you really got it? ok what's the instanteous velocity of s=t^4-6^2+3 at t=3

OpenStudy (anonymous):

Oh and by the way I am a real clown, A Shriner Clown

myininaya (myininaya):

s=t^4-6t^2+3

myininaya (myininaya):

cool. lol

OpenStudy (anonymous):

s=45?

myininaya (myininaya):

what is s' ?

OpenStudy (anonymous):

um, ok im lost ! Help!!

myininaya (myininaya):

we want to find the instanteous velocity of s=t^4-6t^2+3 at t=3 finding velocity means we need to find the derivative of s s' denotes the derivative first let's break this up what is the derivative of t^4? what is the derivative of -6t^2? what is the derivative of 3?

OpenStudy (anonymous):

3t,12t,0

OpenStudy (anonymous):

-12t oops

myininaya (myininaya):

derivative of t^4 is 4*t^3 i used the power rule everything else you got right so 2/3 is passing

OpenStudy (anonymous):

you funny and yes you are right about the first answer , you are making me think faster than normal! and making me nervous! I do not want to come off to you as a complete rock!

myininaya (myininaya):

i'm sorry

myininaya (myininaya):

lol take as much time as you need

OpenStudy (anonymous):

ok, Thanks! Just wish sometimes I was better at this stuff, like you- I love smart math people!

myininaya (myininaya):

clowns are cool too

myininaya (myininaya):

let's put all the terms of s' together so s'=?

OpenStudy (anonymous):

ok, hold, please

myininaya (myininaya):

k

OpenStudy (anonymous):

s=96

myininaya (myininaya):

i don't think thats right and like polak said before its s'(3)=96 tell me what s' is first

myininaya (myininaya):

s'(t)=

OpenStudy (anonymous):

4(3)^3-12+3 Right?

myininaya (myininaya):

s'(3)=4(3)^3-12(3)

myininaya (myininaya):

remember we said the derivative of t^4 is 4t^3 the derivative of -6t^2 is -12t the derivative of 3 is 0

OpenStudy (anonymous):

OK!!!!! 3 is prime right?

myininaya (myininaya):

yes 3 is prime since 3 is a constant, its slope is zero, and since its slope is zero then the derivative of 3 is zero

OpenStudy (anonymous):

we had a- 6t in the original problem? sorry my Bad yes we did, SO SORRY!

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

s= 144+ 3?

OpenStudy (anonymous):

or just 144?

OpenStudy (amistre64):

ok...so whats our question?

myininaya (myininaya):

s=t^4-6t^2+3 find velocity at t=3 i think hes getting mixed up on where things go

OpenStudy (amistre64):

so clown; you say you can find a derivative; right?

OpenStudy (anonymous):

I thought so, guess not

OpenStudy (amistre64):

go ahead and give it a shot, i wanna see what you are doing

OpenStudy (amistre64):

what is the derivative of: t^4 -6^t +3 ??

OpenStudy (anonymous):

Here is my problem..... s=8t^2 -10t +6; t=5

OpenStudy (amistre64):

ok....lets work with 8t^2 -10t +6; what do you think the derivative should look like?

OpenStudy (anonymous):

16t- 10 + 6

OpenStudy (amistre64):

very close; the first and second terms are good; that last terms should be a 0 tho. do you know why?

OpenStudy (anonymous):

because there is no constant?

OpenStudy (amistre64):

does 6 = 6 t^0 ? and if so; how would you derive: 6 t^0?

OpenStudy (anonymous):

Yes , due to the fact of the power 0

OpenStudy (amistre64):

good; then lets rewrite this in terms you can see :) 8t^2 -10t +6t^0; now derive that for me...

OpenStudy (anonymous):

16t-10+6

OpenStudy (amistre64):

16t is good; -10 is excellent. but 6 is wrong, you didnt derive it the same way you did the others ...why not? derive 6t^0. in exactly the same way as you did the others. what do we get?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

that is correct:) so would you agree that our derivative is then: 16t-10 ?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

ds -- = 16t-10 very good; now when t=5, what does this equal? dt

OpenStudy (anonymous):

70!! I got it and you did not leave me at the train depot!!! YEA!!!!!

OpenStudy (amistre64):

yay!!! :)

OpenStudy (anonymous):

Thaank you very much!!!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

Talk to ya soon maybe